The mean number of errors due to a bug occurring in a minute is 0.0001. What is the
probability that no error will occur in 20 minutes
1
Expert's answer
2020-06-04T19:30:57-0400

Here we can use a Poisson Distribution with mean "m=0.0001" in a minute.

"\therefore" for no error in 20 minutes equivalent "m=20(0.0001)"

"=" "0.002"


now Poisson distribution says


"P(X=x) = e\raisebox{0.25em}{-m} * \frac{m^x}{x!}"


therefore in order to find the probability no error will occur in 20 minutes we put "x=0."


"P(X=0)= e\raisebox{0.25em}{-0.002}*\frac{(0.002)^0}{0!}"

"= 0.998*1"

"=0.998"

Answer : 0.998